1)A sales promoter offers a special price for a scanner with a printer. The probabilities that each functions satisfactorily over three years are 0.68 and 0.72 respectively. What is the probability that at the end of the three years:
i) Both scanner and printer will function satisfactorily?
ii) Only the printer will be functioning?
iii) At least one of the two machines will be functioning? State any assumptions made.
-------------------------------------------------------------------------
2)240 students were asked to taste three types of cola in a random order and were not told their identity. The contingency table below shows the counts of preferred colas classified by the subject specialism of students.

Cola A Cola B Cola C

Mathematics 13 23 44

Economics 36 24 10

Statistics 24 8 8

Management 11 28 11


Use These data to investigate the association, if any, between a student's subject specialism and preferred cola-type. Interpret any association which exists, testing at more than one significance level.
--------------------------------------------------------------------------------

3)The times required by three workers to perform an assembly-line task were recorded on five randomly selected occasions. Here are the times, to the nearest minute.



Hank Joseph Susan
8 8 10
10 9 9

9 9 10

11 8 11

10 10 9


Assuming the normality of populations under consideration, at the 5% significance level,
do the data provide sufficient evidence to conclude that a difference exists between the times required by three workers to perform an assembly-line task?

Your kind assistance is highly appreciated
Bahman Mennati








Answer by stanbon(48569)

1)A sales promoter offers a special price for a scanner with a printer. The probabilities that each functions satisfactorily over three years are 0.68 and 0.72 respectively. What is the probability that at the end of the three years:
i) Both scanner and printer will function satisfactorily?
Ans: (0.68*0.72) = 0.4896..
ii) Only the printer will be functioning?
Ans: 0.72*(1-0.68) = 0.2304..

iii) At least one of the two machines will be functioning? State any assumptions made.
Ans: At least one = 1 - P(none)
At least one = 1 - (1-0.72)(1-0.68) = 0.9104
-------------------------------------------------------------------------
2)240 students were asked to taste three types of cola in a random order and were not told their identity. The contingency table below shows the counts of preferred colas classified by the subject specialism of students.

Cola A Cola B Cola C

Mathematics 13 23 44
Economics 36 24 10
Statistics 24 8 8
Management 11 28 11
Use These data to investigate the association, if any, between a student's subject specialism and preferred cola-type. Interpret any association which exists, testing at more than one significance level.
Ho: The row and column factors are independent.
H1: They are dependent
I ran an ANOVA-test on the data and got:
test statistic: F = 0.05486...
p-value: 0.9469
Since the p-value is greater even then 10%, do not reject
Ho at the 10% significance level.
The row and column factors are independent.
---------------------------------------------------------
--------------------------------------------------------------------------------

3)The times required by three workers to perform an assembly-line task were recorded on five randomly selected occasions. Here are the times, to the nearest minute.
Hank Joseph Susan
8 8 10
10 9 9

9 9 10

11 8 11

10 10 9


Assuming the normality of populations under consideration, at the 5% significance level,
do the data provide sufficient evidence to conclude that a difference exists between the times required by three workers to perform an assembly-line task?
Note: The data posted is too confusing to decipher.
Cheers,
Stan H.